Townson Company is making plans for the introduction of a new product, which has a target...

  1. Townson Company is making plans for the introduction of a new product, which has a target selling price of $7 per unit. The following estimates of manufacturing costs have been derived for 6 million units, to be produced during the first year (Points 2.5):                     
  • Direct material: $6,000,000
  • Direct labor: $2,100,000 (at $14 per hour)


Overhead costs have not yet been estimated, but monthly data on total production and overhead for the past 12 months have been analyzed by using least-squares regression. The major overhead cost driver is direct labor hours, with the following results:

  • Fixed overhead cost: $3,200,000
  • Variable overhead cost per unit (coefficient): $2.25

Prepare the company's cost equation (Y = a + bX) to estimate overhead.

Calculate the predicted overhead cost at an activity level of 6,300,000 units.

           What is Townson’s dependent variable in this case?
Answer #1
Given :
Fixed Overhead cost = $    3,200,000
Variable overhead cost per unit = $              2.25
So Total Overhead cost Equation can
be expressed as Y=a+bX
or Y=3,200,000+2.25*X
Where X =no of units
Predicted overhead when X=6,300,000
Y=3,200,000+2.25*6,300,000
Y =$14,175,000
Townson's dependent variable is
total Overhead cost which depends
on the number of units produced
or X